if three times a boy's ages is increased by3, it will equal his sister's age. The sum of their ages now is 23. how old are the boy and his sister?
Let x=boy's age
3x+3=sister's age x+(3x+3)=23

solve for x. solve for 3x+3.
x=________(boys age) 3x+3=_________(sister's age)
verify.

Answer by ikleyn(52776)

.

The equation in your post

    x + (3x+3) = 23

is correct.


Continue to solve it

    x + 3x = 23 - 3

     4x    =   20

      x    =   20/4 = 5.


The boy is 5 years old.

The sister is  3*5+3 = 18 years old.


CHECK.  5 + 18 = 23  the sum of their ages.    ! correct !

Solved, with explanations.

-----------------

If you want to learn on how to solve simple and simplest linear equations and associated word problems,  read the lessons
    - HOW TO solve a linear equation
    - Simple word problems to solve using a single linear equation
in this site.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.

Happy learning  (!)